Circular No. 6493

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

DO DRACONIS J. A. Mattei, AAVSO, reports that this cataclysmic variable is in out-

burst, as indicated by the following magnitude estimates (visual unless otherwise noted): Oct. 15.766 UT, [14.3 (G. Poyner, Birmingham, England); 15.824, [12.4 (M. Gill, Solihull, England); 16.784, 12.7 (Poyner); 16.787, 12.0 (T. Vanmunster, Landen, Belgium; CCD); 16.956, 11.1 (Gill). The last two outbursts of this dwarf nova, which lasted five days each, were in 1988 Mar. (cf. IAUC 4574) and 1990 Sept. (IAUC 5101), when the star reached magnitudes 10.1 and 9.7 at maximum, respectively.

COMET C/1995 O1 (HALE-BOPP) H. Kawakita, Osaka; M. Fujii, Okayama; and J. Watanabe, National

Astronomical Observatory of Japan, report their detection of NH_{2} emission features on Oct. 3.43 UT with an 0.28-m telescope with a spectrograph having resolution ∼ 1.2 nm (slit size 424′′ × 7′′). From observations of [O I] at 630 nm, and based on Haser’s model (using the typical scale lengths of NH_{2}), the production-rate ratio Q(NH_{2})/Q(H_{2}O) = 0.1.

COMET C/1996 Q1 (TABUR) Extension to ephemeris on IAUC 6479 (elements MPC 27882):

1996TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 14 11^{h}38.^{m}23 +55^{o}49.^{′}5 0.459 0.923 67.^{o}2 85.^{o}4 5.5
19 13 14.17 +52 31.1 0.534 0.889 62.7 85.0 5.6
24 14 11.15 +47 18.1 0.630 0.863 59.2 82.0 5.9
29 14 44.86 +42 20.6 0.736 0.846 56.3 77.4 6.1
Nov. 3 15 06.22 +38 03.1 0.846 0.840 53.7 72.1 6.4
8 15 20.76 +34 22.4 0.955 0.844 51.4 66.5 6.7
13 15 31.30 +31 10.7 1.061 0.859 49.3 61.0 7.0
18 15 39.37 +28 22.1 1.162 0.883 47.6 55.8 7.3
23 15 45.83 +25 52.0 1.255 0.915 46.3 51.3 7.6
28 15 51.21 +23 37.4 1.340 0.955 45.4 47.3 7.9
Dec. 3 15 55.85 +21 36.3 1.417 1.001 44.9 44.1 8.3
8 15 59.91 +19 47.1 1.486 1.052 45.0 41.4 8.6
13 16 03.51 +18 08.7 1.545 1.107 45.5 39.4 8.9
1996 October 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green